OpenStudy (anonymous):

Solve the system by using a table. -0.2x - 0.3y = 0.2 0.4 - 0.3y = 1.4

hero (hero):

I think you actually might need two tables for this. One for each equation.

hero (hero):

By the way, shouldn't it be: -0.2x - 0.3y = 0.2 0.4x - 0.3y = 1.4

OpenStudy (anonymous):

thats what i have

OpenStudy (anonymous):

and how would i make two tables for it?

hero (hero):

 Keep finding y values until you have the same y value in both tables.

OpenStudy (anonymous):

ok one more question how do i find the y values for it.

hero (hero):

Plug the x values into the the equation then Isolate and solve for y to get the corresponding y value. For example, plugging x = 0 into the first equation yields: -0.2(0) - 0.3y = 0.2 -0.3y = 0.2 y = -0.2/0.3 y = -2/3

OpenStudy (anonymous):

do just multiply -0.2(1) and thats my answer?